OpenStudy (anonymous):

use the elimination method to find all solutions of the system: y+x^2=4x y+4x=16

OpenStudy (anonymous):

First, solve for one variable ( usually the easiest) \[y+4x = 16\] so \[y = 16-4x\] Then plugin to the other eqaution \[16-4x+x^2=4x\] Hence \[16+x^2 = 8x\] Leaving the quadratic equation \[x^2-8x+16 = 0\] You can solve this by factoring or using the quadratic formula

OpenStudy (anonymous):

Thank you!! I was stuck on the very last part.

OpenStudy (anonymous):

no problem